Classify this conic section and write it in standard form.
-2y^2-24y+72x-432=0

Answer by stanbon(75887)

Classify this conic section and write it in standard form.
-2y^2-24y+72x-432=0
----
Divide thru by -2:
y^2 + 12y -36x + 216 = 0
---------
y^2 + 12y + 36 + 180 = 36x
(y+6)^2 + 180 = 36x
----
x = (1/36)(y+6)^2 + 5
========================
Parabola opening to the right.
======================================
Cheers,
Stan H.